OpenStudy (falco276):

The equation of a circle is shown below. (x -5)2 + (y + 2)2 = 64 The radius of the circle is ________ units.

OpenStudy (falco276):

@Gin_Ichimaru @micahwood50 @Loujoelou @surjithayer @maywadee313 @Zale101 @zepdrix

zepdrix (zepdrix):

\[(x -5)^2 + (y + 2)^2 = r^2\] \[r^2=64, \qquad r=?\]

OpenStudy (zale101):

if r^2=64 then r equals what?

OpenStudy (falco276):

I actually don't know. :(

zepdrix (zepdrix):

So we have r^2=64. To solve for r, take the square root of both sides. r=sqrt(64). Use a calculator if you aren't sure about this one! :)

OpenStudy (falco276):

erm, okay.

OpenStudy (falco276):

What do you mean by both sides?

OpenStudy (falco276):

I got 4096. :(

zepdrix (zepdrix):

\[\Large r^2=64\]In the land of math, we can manipulate both sides and the equation stays balanced. So we'll take the square root of r^2 and the square root of 64.\[\Large \sqrt{r^2}=\sqrt{64}\]Squareroot and square are inverse operations of one another, they'll undo each other in this instance.\[\Large r=\sqrt{64}\] Hmm 4096? It sounds like you `squared` your 64. You want to `square root` your 64.

OpenStudy (falco276):

How? Did I happen to Multiply 64 by 64? oops, my mistake. *blushing*

OpenStudy (falco276):

Anyone? :(

zepdrix (zepdrix):

Yes you multiplied 64 and 64 :) You want to instead use your calculators \(\large \sqrt{}\quad\)button.

OpenStudy (falco276):

okay. 8

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (falco276):

Aww, thanks. Imma medal you! :D